OpenStudy (anonymous):

@jdoe0001 @texaschic101 what is the fraction 1 588/756 in simplest form?

OpenStudy (anonymous):

16/9

OpenStudy (anonymous):

and as a mixed number would be 1 and 7/9?

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

1 588/756 = (1 * 756 + 588) / 756 = (756 + 588) / 756 = 1344/756 = (divide numerator and denominator by 84) = 16/9 or 1 7/9

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you are welcome